OpenStudy (anonymous):

Find the Taylor polynomials of degree n approximating 4/(2−2x) for x near 0: for n = 3 for n = 5 for n = 7

ganeshie8 (ganeshie8):

for x near 0, taylor series is gives the maclaurin series

ganeshie8 (ganeshie8):

\( f(x) = f(0) + xf'(0) + \dfrac{x^2}{2!}f''(0) + \dfrac{x^3}{3!}f'''(0) + ... \)

ganeshie8 (ganeshie8):

maybe start by finding the \(n\)th derivative of given function

OpenStudy (anonymous):

Alright. I'll try that in the morning. It's roughly 4 am here, so I'm getting some sleep. Thanks!

ganeshie8 (ganeshie8):

have good sleep :)

OpenStudy (amistre64):

the long division creates the poly ... 2+2x+2x^2+.... seems to be filling out rather simply ----------------------- 1-x ) 2 -(2-2x) ------ 2x -(2x-2x^2) ----------- 2x^2